# MDA: An Interpretable Multi-Modal Fusion with Missing Modalities and Intrinsic Noise

### v2

@@ Introduction
Fig. 3 illustrates the average modal attention of the proposed method for different disease categories. Firstly, the model exhibits notable specificity in modal attention for different disease categories. It is important to note that no constraints were imposed during model training regarding which modalities should be attended to for the recognition of different diseases. We attribute this specificity to the proposed modality weight adaptation module working in conjunction with multi-disease classification. Furthermore, we analyzed the alignment between the model’s specificity in modal attention for different disease categories and clinical priors. According to a study on gastrointestinal diseases by Jacobson et al. (Jacobson et al. 2023), using WLE alone to diagnose lipomas achieves a specificity of 99% in clinical practice. Similar findings hold true for NET. Conversely, GIST and EP rely more heavily on EUS. This alignment between the MDA weight and clinical expertise is highly consistent (Fig. 3), providing interpretability of modal attention for each disease category.
@@
Table 6: Variations in adaptive weights for different modalities in the absence of certain modalities and learning with intrinsic noise.
Disease Modal Baseline w/o EUS w/o WLE w/o Report
@@
Micro-visualization of model adaptation to missing modalities Table 6 demonstrates the variations in attention weights for three modalities before and after incorporating missing modalities and intrinsic noise in the network. The findings indicate that the proposed model is highly sensitive to these changes, as evidenced by the rapid and substantial shifts in attention weights across modalities when dealing with missing modalities. When EUS is missing, a significant decrease in EUS weights is first observed (0.34 to 0.05, 0.53 to 0.11), which is also observed when WLE is missing. When WLE is missing, attention weights almost entirely shift towards the report modality, which is expected since the surface information of the tumor, reflected in WLE, cannot be found in EUS. When WLE is absent, the model cannot learn relevant knowledge solely from the EUS modality. However, the report modality contains descriptions of WLE and can serve as a substitute for the missing image modality. Hence, the weights assigned to the report modality increase. Similarly, in the absence of EUS, the emphasis shifts more towards the reporting modality rather than WLE modality. In the scenario where the report modality is missing, the redistribution of the reduced weight is approximately evenly allocated to both WLE and EUS to compensate for the absence. Due to the lack of information provided by the missing modality, experiments indicate that when a missing modality is present (i.e., when the modality does not provide information), the MDA module can effectively reduce the impact of the missing modality on the final results through weight redistribution.
Conclusion
This paper introduces MDA, which employs a continuous attention mechanism to capture the linear correlations between multiple modalities. MDA exhibits the capability to effectively handle multi-modal information, even when dealing with missing modalities and intrinsic noise, eliminating the need for separate solutions for each scenario. To the best of our knowledge, This work is the first to explore intrinsic noise in diagnostic texts and MDA can withstand this noise interference. We have achieved SOTA results across various tasks and datasets, and have provided novel insights into the interpretability of modalities.
The interpretability of MDA in medical diagnosis is comprehensively analyzed. At the macroscopic level, we investigate the attention specificity of MDA towards different disease categories, demonstrating its alignment with clinical experience. At the microscopic level, we examine the significant changes in attention exhibited by MDA when handling missing modalities. The experimental results provide evidence of the efficacy of MDA in improving multi-modal fusion and its robustness in the presence of missing modalities or intrinsic noise. Additionally, we have made available a multi-modal dataset of gastrointestinal stromal tumors for future research purposes.

### v3

@@ 4.3 Interpretability analysis
MDA exhibits notable specificity in modal attention for different disease categories. It is important to note that no constraints were imposed during model training regarding which modalities should be attended to for the recognition of different diseases. We attribute this specificity to the proposed modality weight adaptation module working in conjunction with multi-disease classification. Furthermore, we analyzed the alignment between the model’s specificity in modal attention for different disease categories and clinical priors. In clinical practice, multi-modal imaging techniques are frequently employed to gather comprehensive diagnostic information. For instance, the combination of EUS, WLE, CT, and MRI is utilized for the diagnosis of gastrointestinal diseases, with each imaging modality being sensitive to different features. According to a study on gastrointestinal diseases [36, 5, 14], using WLE alone to diagnose lipomas achieves a specificity of 99% in clinical practice. Similar findings hold for NET. Conversely, GIST and EP rely more heavily on EUS  [14]. This alignment between the MDA weights and clinical expertise is highly consistent. The findings can corroborate existing clinical experience and serve as a reference for modalities lacking such experience, which is precisely what is needed in the current field of multi-modal medical algorithms: reinforcing the role of artificial intelligence in establishing the gold standard for diagnostic use of multi-modality.
@@
Table 6: Variations in MDA weights in missing modalities and learning with intrinsic noise.
Disease Modal Baseline w/o EUS w/o WLE w/o Rep.
@@
Micro-visualization of model adaptation to missing modalities Table 6 demonstrates the variations in MDA weights for three modalities before and after incorporating missing modalities in the network. The findings indicate that the proposed model is highly sensitive to these changes, as evidenced by the rapid and substantial shifts in modal-domain MDA weights across modalities when dealing with missing modalities. When EUS is missing, a significant decrease in EUS weights is first observed (0.34 to 0.05, 0.53 to 0.11), which is also observed when WLE is missing. When WLE is missing, MDA weights almost entirely shift towards the report modality, which is expected since the surface information of the tumor, reflected in WLE, cannot be found in EUS. When WLE is absent, the model cannot learn relevant knowledge solely from the EUS modality. However, the report modality contains descriptions of WLE and can serve as a substitute for the missing image modality. Hence, the weights assigned to the report modality increase. Similarly, in the absence of EUS, the emphasis shifts more toward the reporting modality rather than the WLE modality. In the scenario where the report modality is missing, the redistribution of the reduced weight is approximately evenly allocated to both WLE and EUS to compensate for the absence. Due to the lack of information provided by the missing modality, experiments indicate that when a missing modality is present (i.e. when the modality does not provide information), the MDA module can effectively reduce the impact of the missing modality on the final results through weight redistribution.
5 Conclusion
The proposed MDA represents a pivotal advancement in multi-modal fusion, addressing the core challenges of missing modalities, intrinsic noise, and interpretability. MDA underscores the significance of a unified technical framework, which successfully constructs linear attention relationships between various modalities, permits adaptive control over modality weights, and scales to an arbitrary number of modalities with lower complexity.
This study pioneers the exploration of interpretability for key features and gold standard modalities in computer-aided diagnosis through the visualization of macroscopic multi-modal MDA maps. Its objective is to enhance the contribution of artificial intelligence to the standardization of diagnostic criteria. Furthermore, by visualizing the attention shifts in the internal correlations across multiple modalities at a microscopic level, the research imbues the model with self-explanatory capabilities.